OpenStudy (anonymous):

Solve this problem. A student needs at least 80% of all points on the tests in a class to get a B. There are three 100-point tests and a 250-point final. The student's test scores are 72%, 85% and 78% for the 100 point tests. What is the least score the student can make on the final and still get a B?

OpenStudy (curry):

a 205

OpenStudy (curry):

dont forget medal

OpenStudy (lalaly):

35 + x = 320 x = 320-235 = 85 x = 85% 250 Question Test .85*250 = 212.5 Students needs a score of at least 213 on last test

OpenStudy (lalaly):

the first number is 235 *** sory

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

if y varies directly as x^2, find k if y = 27 and x = 3